# Measurement induced phase transition in a quantum Ising model by periodic measurements

### v1

@@ IV GENERALIZED GEOMETRIC MEASURE
Figure 6: Behavior of multipartite entanglement with time-step $n$ and $\tau$ : In panel (a), we have plotted $S_{G}(n)$ vs $n$ at field strength $h=\frac{1}{2}$ for different $\tau$ values. Here, the $\tau$ values are taken from $\tau=0.1,0.2,.....,1.0$. A gap between the curves corresponding to the $\tau$ values is detected. Next, in panel (b), $\mathcal{G}_{n}$ vs $n$ is plotted. Here, $\mathcal{G}_{n}$ is the cumulative average of the sequence $\bigl{\{}S_{G}(1),S_{G}(2),S_{G}(3),S_{G}(4)...S_{G}(n)\bigl{\}}$. In panel (c) we have plotted $\frac{d\mathcal{G}_{n}}{d\tau}$ vs $n$ plot, where the peak value occurs around $\tau=\tau_{c}=0.2$. It depicts the transition that happens near $\tau_{c}=0.2$.
V Recursion relation for survival probability
It has been shown earlier Dhar and Dasgupta (2016) that the transition when the measurement is complementary to the question (3), namely,
Is the magnetic moment (per site) $M_{z}\neq 1$? (11)
can be treated analytically for any arbitrary system size. It was proved that the ratio of two successive first detection probabilities is given by
$\frac{p_{n}}{p_{n-1}}=e^{-\alpha L\tau^{2}}$ (12)
with
$\alpha(h,\tau)=-\frac{1}{L\tau^{2}}\sum_{k=0}^{\pi}\,\log\left[1-\left(\frac{2\sin k\,\sin(\lambda_{k}\tau)}{\lambda_{k}}\right)^{2}\right]$ (13)
The quantity $\alpha$ shows a change of slope, indicating a criticality, as a function of $h$ and $\tau$. However, as is evident from Eq. (12) the first detection probability is finite only for finite size and vanishes in the thermodynamic limit.
@@ V Recursion relation for survival probability
For the present measurement (3), we could not derive a closed-form expression for the survival probability, but we could derive a recursion relation that enables one to compute the probability $R_{n}$ for large size. To this end, we introduce the quantities
@@
Using the recursion relations, one can find the survival probability after $n$ measurements as
Figure 7: Survival probability at large system sizes : In panel (a), we have plotted the derivative of $H$ vs. $\tau$ graph for $L=100,200,300,500,1000$, where $H$ is denoted as the height of the plateau region for a given $\tau$ obtained by using Eq. (24). It is apparent that $\tau_{c}$ moves towards 00 as $L$ increases. In panel (b), we have plotted $\frac{dH}{d\sigma}$ vs. $\sigma$ for various system sizes, that is, $L=100,200,300,500,1000$ at field $h=\frac{1}{2}$ where $\sigma=\tau\sqrt{L}$. All the curves have a peak around $\sigma_{c}=1$
@@
This expression is valid for any Hamiltonian, provided $f_{n}$ are known. For the specific case of transverse Ising Hamiltonian, one can derive an expression for $f_{n}$ in closed form for any even value of $L$, by using the exact solution Damski and Rams (2013). One can transform spin variables $s_{j}$ to fermion variables $a_{j}$ and perform Fourier transformation to fermion variables $a_{k}$, to get $\mathcal{H}$ as a Kronecker sum of commuting Hamiltonians $\mathcal{H}_{k}$ :
$\mathcal{H}$ $=$ $\sum_{k=0}^{\pi}\mathcal{H}_{k},\;\;\mathcal{H}_{k}=-2i\sin k\left[a_{k}^{\dagger}a_{-k}^{\dagger}+a_{k}a_{-k}\right]$ (25)
$-2(\Gamma+\cos k)\left[a_{k}^{\dagger}a_{k}+a_{-k}^{\dagger}a_{-k}-1\right]$
@@
where $k=(2n+1)\pi/L$ with $n=0,1,2,\cdots,L/2-1$. Each of these $\mathcal{H}_{k}$’s can be described by four basis states namely, $|00\rangle_{k}$, $|11\rangle_{k}$, $|10\rangle_{k}$ and $|01\rangle_{k}$, where the numbers in each basis signify the occupation status of the fermions having momenta $+k$ and $-k$ respectively. (We consider the even-occupation states only.) The ground state and excited state of $\mathcal{H}_{k}$ are
@@
[…]
$e^{-i\mathcal{H}_{k}n\tau}|11\rangle_{k}$ $=$ $\left[\cos(\lambda_{k}n\tau)+i\sin(\lambda_{k}n\tau)\cos(2\theta_{k})\right]|11\rangle_{k}$ (27)
$-\sin(2\theta_{k})\sin(\lambda_{k}\tau)|00\rangle_{k}$
@@
In principle, one should be able to calculate the survival probability using Eqs. (20, 24, 28), but due to some precision problems we could not go beyond $L=1000$ and $n=10000$.
As for small size, the survival probability shows a plateau region and the derivative of the height $H$ of the plateau with respect to $\tau$ shows a peak at a critical point $\tau=\tau_{c}$. (Fig. 7). It is found that if one uses instead of $\tau$ a scaled variable, $\sigma=\tau\sqrt{L}$ then the plot of $dH/d\sigma$ vs $\sigma$ coincide for different values of $L$. This indicates that the position of the peaks $\tau_{c}$ varies as $1/\sqrt{L}$ and vanishes for large $L$. We shall derive below this scaling law in the limit of small $\tau$ and small $n\tau$.
To this end, we first write $f_{n}$ in Eq. (28) as
@@
and hence the argument of each summand in the expression for $R_{n}$ in Eq. (24) is zero. Thus, the main contribution to $R_{n}$ comes from $\rho_{n}$, which, in the limit of small $n\tau$ becomes
@@
This indicates that the parameters $\tau$ and $L$ may be expected to occur as the combination $\tau^{2}L$ in the value of survival probability.
VI Summary
We present the observation of measurement-induced phase transition in a quantum spin model with projective non-random measurement. We start with a chain of $L$ Ising spins all oriented in the +Z direction. In each time step, it evolves unitarily under the transverse Ising Hamiltonian for time $\tau$, and then a measurement is performed that gives a binary answer to the question, “Is the magnetic moment per site equal to 1?”. After $n$ time-steps, the probability that the answer is “no” in each case is termed as the survival probability $R_{n}$. We have derived a recursion relation, using which we could compute $R_{n}$ for size up to $L=1000$. The plot of $R_{n}$ vs. $n$ shows a plateau region at small $n$. The key parameter is the rate at which the height (say $H$) of this plateau changes with $\tau$. The plot of $\frac{dH}{d\tau}$ shows a peak at some value of $\tau$, which we identify as the critical point $\tau_{c}$, since the bipartite and multipartite entanglement undergoes an area-law to volume-law transition there. We prove analytically that, for small $\tau$, the critical point $\tau_{c}$ scales as $1/\sqrt{L}$. On using the scaled variable $\sigma=\tau\sqrt{L}$, one finds that the plots of $\frac{dH}{d\sigma}$ vs $\sigma$ for different system sizes are coincident, and as $L$ increases, the peak in the plot of $\frac{dH}{d\tau}$ rises in height and shifts towards zero. Hence, the transition is only observable for finite $L$. Interestingly, the survival probability decays logarithmically with time in some cases. All these cases correspond to the phase for which the ground state of the Hamiltonian is disordered, indicating that this transition is dependent on the presence of order in the ground state of the Hamiltonian. However, there are some values of the external field for which the ground state is disordered, but the logarithmic decay is not present. Also, one must note that all the results discussed above are for one particular choice of the initial state.
Acknowledgements
PN acknowledges UGC, India for financial support (Ref. No. 191620072523) and Harish Chandra Research Institute for access to their infrastructure.
Appendix A
Here, in appendix A, we do the same analysis as the main paper but at field strength $h=\frac{3}{2}$. Firstly, in FIG. A.1, we have shown the $R_{n}$ vs. $n$ plot at $h=\frac{3}{2}$. Here, the notation $R_{n}$ and $n$ are consistent with the main paper. The same feature of the curves corresponding to the $\tau$ value $0.1$ and $0.2$ is captured.
Figure A.1: $R_{n}$ vs $n$: Here, we have plotted $R_{n}$ vs. $n$ graph corresponding to each $\tau$ value for system size $L=28$ at field $h=\frac{3}{2}$.
Figure A.2: $\frac{dH}{d\tau}$ vs $\tau$ plot. We have plotted $\frac{dH}{d\tau}$ vs. $\tau$ for different system sizes, that is, $L=20,22,24,26,28$ at field $h=\frac{3}{2}$. The curves precisely illustrates the $\tau$ values at which this transition happens. The transition point in this case is represented by the highest values of the curves.
We again plot the derivative of the height of the plateaus from FIG. A.1 with respect to $\tau$. Then, we see the behavior of $\frac{dH}{d\tau}$ vs $\tau$ in FIG. A.2 for various system sizes, that is, $L=20,22,24,26,28$. Here, the peaks of the curves are identified as the critical point. The transition point moves towards a lower value of $\tau$ as it is seen in the study for $h=\frac{1}{2}$.
@@ Acknowledgements
Figure A.3: In panel (a), entanglement entropy for the whole system in $L/4:\text{rest}$ bipartition ($S(L/4)$) is plotted for different $\tau$ values with respect to $L$, which is the system size. We have taken $n=10$ and calculated entanglement and see an area law to volume law transition. $\tau$ values increase as the color of the curve becomes deeper. In panel (b) we have plotted $S(L/4,\tau)$ for different values of $L$ with respect to $\tau$. In panel (c) we have shown the scaling collapse of the entanglement curves occurs around $\tau_{c}=0.1$ with the exponent $0.4$.
Figure A.4: In panel (a), the plot of $S_{G}(n)$ vs $n$ at field strength $h=\frac{3}{2}$ for different $\tau$ values. Here, the $\tau$ values are taken from $\tau=0.1,0.2,.....,1.0$. A gap between the curves corresponding to the $\tau=0.1$ and $\tau=0.2$ values is detected. Next in panel (b), $\mathcal{G}_{n}$ vs $n$ is plotted at the same field strength, that is, $h=\frac{3}{2}$. In panel (c) we have plotted $\frac{d\mathcal{G}_{n}}{d\tau}$ vs $n$ for $L=26$. plot, which shows the transition point around $\tau_{c}=0.1$.
After that we calculate bipartite entanglement and show from FIG. 3(a) the area to volume law transition occurs apparently at $\tau_{c}=0.1$. at field strength $h=\frac{3}{2}$. This area-to-volume law transition is also detected from the collapse of curves of different system sizes $L$ at $\tau<\tau_{c}=0.1.$. It is shown at 3(b). At last in FIG. 3(c) we show the scaling collapse at $\tau_{c}=0.1$ and $\mu=0.4$.
Then we also analyze stochastically attained GGM (SAG) at the field strength $h=\frac{3}{2}$. In FIG. 4(a) we have plotted $S_{G}(n)$ vs $n$ for ten $\tau$ values from $0.1$ to $1.0$ with an interval $0.1$. and in FIG. 4(b)  $\mathcal{G}_{n}$ vs $n$ plot is depicted for the same set of $\tau$ values as shown in FIG. 4(a). Here also, the notations $S_{G}(n)$ and $\mathcal{G}_{n}$ are consistent with the main text. At last in FIG. 4(c) from $\frac{d\mathcal{G}_{n}}{d\tau}$ curves for system size $L=26$, the transition occurs around the point $\tau_{c}=0.1$.
@@
Figure A.5: In panel (a), we have plotted $\frac{dH}{d\tau}$ vs. $\tau$ for $L=100,200,300,500,1000$ at the field strength $h=\frac{3}{2}$. Even at $h=\frac{3}{2}$, $\tau_{c}$ also moves towards 00 as $L$ increases. In panel (b), we have plotted $\frac{dH}{d\sigma}$ vs. $\sigma$ at the same system sizes and field strength as FIG. 5(a), where $\sigma=\tau\sqrt{L}$. All the curves have a peak around $\sigma_{c}=1$.
At last, in FIG. A.5 we do some large system size studies at field strength at $h=\frac{3}{2}$. In FIG. 5(a) we show the variation of $\frac{dH}{d\tau}$ with $\tau$ at field strength $h=\frac{3}{2}$ for $L=100,200,300,500,1000$. The highest value of $\frac{dH}{d\tau}$ vs $\tau$ curve shows the deviation of the transition point $\tau_{c}$ towards 00 with increase of the system size. In FIG. 5(b) we show the $\frac{dH}{d\sigma}$ vs. $\sigma$ plot. Here the quantity $\sigma=\tau\sqrt{L}$. As all the curves corresponding to different system sizes have peak values at $\sigma_{c}=1$, it shows that $\tau_{c}\sim\frac{1}{\sqrt{L}}$.

### v3

@@ Abstract
Figure 4: Variation of multipartite entanglement with time-step $n$ and $\tau$ for $h=\frac{1}{2}$: In panel (a), we have plotted $S_{G}(n)$ vs $n$ for different $\tau$ values. The curves for $\tau>0.2$ are nearly coincident, but are separated from the one for $\tau=0.2$. In panel (b), $\mathcal{G}_{n}$ vs $n$ is plotted at $L=26$, where $\mathcal{G}_{n}$ is the cumulative average, namely the average of the sequence $\bigl\{S_{G}(1),S_{G}(2),S_{G}(3),S_{G}(4)...S_{G}(n)\bigl\}$. In panel (c) we have plotted $d\mathcal{G}_{n}/d\tau$ vs $\tau$, where the peak occurs around $\tau=0.2$, indicating that $\tau_{c}=0.2$.
Generalized Geometric Measure:
For quantifying multipartite entanglement, we utilize the measure referred to as the generalized geometric entanglement (GGM) Sen (De); Das et al. (2016); Barnum and Linden (2001); Plenio and V. (2001); Wei and Goldbart (2003); Osterloh and Siewert (2005); Yu and Song (2006); Orús (2008). The GGM of a pure $L$-partite state, $|\psi_{n}\rangle$ is defined as
[…]
$G_{n}=1-\max_{\ell}\left\{\lambda_{\ell,L-\ell}^{2}\right\}_{l}$ (9)
where $\lambda_{\ell,L-\ell}$ is the maximal Schmidt coefficient in the bipartite split $\ell:L$-$\ell$ of our chain. Since we calculate the GGM for the state $|\psi_{n}\rangle$ after the $n$th measurement, we plot the stochastically attained GGM (SAG), defined as
$S_{G}(n)=R_{n}G_{n}$
The plot of $S_{G}(n)$ with $n$ at field strength $h=1/2$ as depicted in the panel (a) of Fig. S.9 for equispaced $\tau$ values shows the presence of plateaus (just as in the case of survival probability) in spite of strong fluctuations. Also, the plateaus are found to be more separated (vertically) near $\tau_{c}$. To avoid fluctuations, we perform cumulative averaging of the data for each $\tau$ value (panel (b) of Fig. S.9). The cumulative average for a sequence, $\bigl\{x_{1},x_{2},x_{3},\ldots,x_{n}\bigl\}$, is defined as $\mathcal{M}_{i}=\frac{1}{i}\sum_{k=1}^{i}x_{k}$ for $i=1,2,\ldots n$. Using the cumulative average, we plot (panel (c) of Fig. S.9) the derivative of the height of the plateau region and find a peak at $\tau_{c}=0.2$, which we identify as the transition point.
Thus, for sizes $L\leq 28$, the three quantities, namely survival probability, bipartite entanglement entropy, and genuine multipartite entanglement, show a measurement-induced transition at a common value of $\tau_{c}$ around $0.2$ at field strength $h=1/2$.
Recursion Relation for Survival Probability:
@@
Although we could not derive a closed-form expression for the survival probability, it is possible to derive a recursion relation that enables one to compute the probability $R_{n}$ for large system-size. To this end, we introduce the quantities
@@
(Similar recursion relations has been discussed extensively Friedman et al. (2017); Liu et al. (2024) in a somewhat different context.) Using the recursion relations, one can find the survival probability after $n$ measurements as
@@
This expression is valid for any Hamiltonian, provided $f_{n}$ are known. For the specific case of transverse Ising Hamiltonian, one can derive an expression for $f_{n}$ in closed form for any even value of $L$, by using the exact solution Damski and Rams (2013) (See Supplementary Material)
where $\lambda_{k}$ and $\theta_{k}$ are defined by
$\lambda_{k}=2\sqrt{h^{2}+1+2h\cos k},\;\;\;e^{2i\theta_{k}}=\frac{2(h+e^{-ik})}{\lambda_{k}}$
In principle, one should be able to calculate the survival probability for any $L$ and $n$ using Eqs. (16, 20, 25), but due to some precision problems we could not go beyond $L=1000$ and $n=10000$.
As for small size, the survival probability shows a plateau region and the derivative of the height $H$ of the plateau with respect to $\tau$ shows a peak at a critical point $\tau=\tau_{c}$. (Fig. 5). It is found that if one uses instead of $\tau$ a scaled variable, $\sigma=\tau\sqrt{L}$ then the plot of $dH/d\sigma$ vs $\sigma$ coincide for different values of $L$. This indicates that the position of the peaks $\tau_{c}$ varies as $1/\sqrt{L}$ and vanishes for large $L$. One can derive analytically this scaling law in the limit of small $\tau$ and small $n\tau$ (see Supplementary  I). Thus, the transition exists only for finite $L$ and vanishes in the thermodynamic limit. We would like to add that scaling of $\tau_{c}$ on the basis of small-size data does not give this $1/\sqrt{L}$ scaling; rather, $|\tau_{c}(L)-\tau_{c}(\infty)|$ scales as $\log L$. The detailed description is provided in supplementary materials I.
Figure 5: Survival probability at large system sizes. (a) : Plot of $dH/d\tau$ vs $\tau$, where it is clear that the peak at $\tau_{c}$ moves towards 0 as $L$ increases. (b): Plot of $dH/d\sigma$ vs. $\sigma$ for various system sizes at field $h=\frac{1}{2}$ where $\sigma=\tau\sqrt{L}$. All the curves coincide and show a peak at $\sigma=1$, indicating the scaling $\tau_{c}\sim 1/\sqrt{L}$.
There are some comments (1) For quantum circuits, the transition arises from a competition between the measurement process, which decreases the entanglement and the unitary evolution, which increases the entanglement. In contrast, for our case the transition arises in spite of the fact that in some cases the measurement increases the entanglement, while in some cases it decreases (Fig. S.9 of the supplementary materials I). (2) The end of the plateau region in the $R_{n}$ vs $n$ plot is accompanied by a sudden dip in the entanglement (Fig. S.10 of the supplementary materials I). Thus, the presence of the plateau in the $R_{n}$ vs $n$ plot seems to be related to the entanglement entropy. (3) As mentioned earlier, the critical value $\tau_{c}$ decreases with system size. Numerical results show that it also depends on the transverse field $h$ (see Supplementary Materials for details). (4) Normally, the MIPT observed in quantum circuits is independent of the parameters of the Hamiltonian. But in our case, we observe that the survival probability at long time shows a logarithmic decay for $h=3/2$, when the ground state of Hamiltonian is disordered (see Supplementary Materials). (5) One must note that all the results discussed above are for the initial state where all the spins are pointed up in the Z direction. We also see for the initial Neel state and field strength $h=\frac{1}{2}$, there exists phase transition, although for initially oriented spins along the $X$ direction, the area-to-volume law transition is observed by bipartite entanglement but is not captured by survival probability.
Discussion:
We present the observation of measurement-induced phase transition in a quantum spin model with projective non-random measurement. We start with a chain of $L$ Ising spins all oriented in the +Z direction. In each time step, it evolves unitarily under the transverse Ising Hamiltonian for time $\tau$, and then a measurement is performed that gives a binary answer to the question, “Is the magnetic moment per site equal to 1?”. After $n$ such time-steps, the probability that the answer is “no” in each case is termed as the survival probability $R_{n}$. We have derived a recursion relation, using which we could compute $R_{n}$ for size up to $L=1000$. The plot of $R_{n}$ vs. $n$ shows a plateau region at small $n$. The key parameter is the rate at which the height (say $H$) of this plateau changes with $\tau$. The plot of $\frac{dH}{d\tau}$ shows a peak at some value of $\tau$, which we identify as the critical point $\tau_{c}$, since the bipartite and multipartite entanglement undergoes an area-law to volume-law transition there for small size. We prove analytically that, for small $\tau$, the critical point $\tau_{c}$ scales as $1/\sqrt{L}$. On using the scaled variable $\sigma=\tau\sqrt{L}$, one finds that the plots of $\frac{dH}{d\sigma}$ vs $\sigma$ for different system sizes are coincident, and as $L$ increases, the peak in the plot of $\frac{dH}{d\tau}$ rises in height and shifts towards zero. Hence, the transition is only observable for finite $L$. As mentioned in the Introduction, MIPT has been observed numerically in a variety of systems and in different contexts. It will be interesting to explore size-dependence of MIPT in these situations.
Acknowledgements:
PN acknowledges UGC, India for financial support (Ref. No. 191620072523) and Harish Chandra Research Institute for access to their infrastructure. SDG acknowledges Eli Barkai for insightful discussions in the conference titled Quantum Trajectories held in ICTS, India.
I Supplemental Material
Derivation for $f_{n}$ in Eq. (21):
As mentioned in the text, for transverse Ising Hamiltonian, one can derive an expression for $f_{n}$ in closed form for any even value of $L$, by using the known analytic solution Damski and Rams (2013). One can transform spin variables $s_{j}$ to fermion variables $a_{j}$ and perform Fourier transformation to fermion variables $a_{k}$, to get $\mathcal{H}$ as a Kronecker sum of commuting Hamiltonians $\mathcal{H}_{k}$ :
$\mathcal{H}$ $=$ $\sum_{k=0}^{\pi}\mathcal{H}_{k},\;\;\mathcal{H}_{k}=-2i\sin k\left[a_{k}^{\dagger}a_{-k}^{\dagger}+a_{k}a_{-k}\right]-2(h+\cos k)\left[a_{k}^{\dagger}a_{k}+a_{-k}^{\dagger}a_{-k}-1\right]$ (22)
@@ I Supplemental Material
where $k=(2n+1)\pi/L$ with $n=0,1,2,\cdots,L/2-1$. Each of these $\mathcal{H}_{k}$’s can be described by four basis states namely, $|00\rangle_{k}$, $|11\rangle_{k}$, $|10\rangle_{k}$ and $|01\rangle_{k}$, where the numbers in each basis signify the occupation status of the fermions having momenta +k+k and k-k respectively. (We consider the even-occupation states only.) The ground state and excited state of $\mathcal{H}_{k}$ are
@@
with eigenvalues $\mp\lambda_{k}=\mp 2\sqrt{h^{2}+1+2h\cos k}$ respectively. Here, $\theta_{k}$ is defined by $e^{2i\theta_{k}}=2(h+e^{-ik})/\lambda_{k}$. To proceed further, we note that since $\sum_{j}s^{z}_{j}=\sum_{k}(2a_{k}^{\dagger}a_{k}-1)$, the state $|0\;0\;0\cdots\rangle$ corresponds to the one with all $k$-modes (positive and negative) occupied. Using Eq. (23), one then obtains
$e^{-i\mathcal{H}_{k}n\tau}|11\rangle_{k}$ $=$ $\left[\cos(\lambda_{k}n\tau)+i\sin(\lambda_{k}n\tau)\cos(2\theta_{k})\right]|11\rangle_{k}-\sin(2\theta_{k})\sin(\lambda_{k}\tau)|00\rangle_{k}$ (24)
@@
Derivation of the scaling law $\tau_{c}\sim 1/\sqrt{L}$:
We first write $f_{n}$ in Eq. (25) as
@@
and hence the argument of each summand in the expression for $R_{n}$ in Eq. (20) (of main text) is zero. Thus, the main contribution to $R_{n}$ comes from $\rho_{n}$, which, in the limit of small $n\tau$ becomes
@@
This indicates that the parameters $\tau$ and $L$ may be expected to occur as the combination $\tau^{2}L$ in the value of survival probability, leading to the scaling law $\tau\sim 1/\sqrt{L}$.
Dependence of the critical point on system-size and external field:
As discussed in the main text, the location of the critical point $\tau_{c}$ depends on the size of the system $L$ and the transverse field $h$ as shown in FIG. S.6
Figure S.1: Critical point $\tau_{c}$ as obtained from the survival probability (see text) as a function of the system size $L$ and field strength $h$.
Dependence of Survival probability decay on transverse field strength:
In FIG (S.2), the survival probability is studied at large $n$ at $h=\frac{3}{2}$ for different $\tau$ values. We can see that $R_{n}$ decays in logarithmic fashion at large $n$. This kind of logarithmic decay is not present for $h=1/2$. The ground state of transverse Ising Hamiltonian is known to be ferromagnetic for $h<1$ and paramagnetic for $h>1$. Thus, although the MIPT in quantum circuits is independent of the parameters of the Hamiltonian, in our case the observable (namely, the survival probability of the initial state) can detect the ground state of the Hamiltonian. However, we must mention that when $h$ is too high, such decay is not present even for $h>1$.
Figure S.2: Survival probability $R_{n}$ vs time-step $n$ for (a) $h=1/2$ and (b) $h=3/2$ at $L=50$. Note the log-scale along X axis and the linear nature of the curve in panel (b).
Results for $h=3/2$:
Survival probability, bipartite entanglement entropy, and multipartite entanglement (stochastically attained GGM) are plotted for $h=3/2$ in Figs. S.3, S.4, S.5 respectively. For large size, the plot of $dH/d\tau$ vs $\tau$ and $dH/d\sigma$ vs $\sigma$ is presented in Fig. S.7
Figure S.3: Survival probability vs time-step and $\frac{dH}{d\tau}$ vs $\tau$ plot. in panel (a) Survival probability vs time-step for different values of $\tau$ at $h=3/2$, $L=28$. In the main figure, we consider $\tau$ values ranging from $\tau=0.1$ to $\tau=0.5$, taken at intervals of $0.1$. At the same time the inset shows same $R_{n}$ vs. $n$ plots corresponding to system size $L=28$ for $\tau$ values ranging from $\tau=0.1$ to $\tau=1.0$, in intervals of of $0.1$. The color bar represents the same $\tau$ values in both the main figure and the inset. In, (b) $\frac{dH}{d\tau}$ vs $\tau$ plot for different system sizes at field $h=\frac{3}{2}$. The transition point is identified as the peak in each curve.
@@
Figure S.4: Bipartite entanglement entropy for $L/4$ : rest bipartition ($S(L/4)$) for $h=3/2$ and $n=10$. In panel (a), entanglement entropy is plotted for different $\tau$ values with respect to the system size $L$. The $\tau$ values increase as the color of the curves becomes deeper. In this context, we have considered values of $\tau$ ranging from $0.06$ to $0.30$, with an increment of $0.02$ between successive $\tau$ values. Now, among all the $\tau$ values, i.e for $\tau=0.06,0.08,0.10,0.12,0.14,....,0.40$, the curves are almost flat bellow $\tau=0.10$ and above $\tau=0.10$, a almost linear increment of entanglement starts to arise. It is apparent that there is an area law to volume law transitionaround $\tau_{c}=0.20$. The transition between area-law and volume-law is clear about the transition point $\tau_{c}=0.1$. It is apparent that there is an area law to volume law transition. In panel (b) we plot $S(L/4,\tau)$ for different values of $L$ with respect to $\tau$. In panel (c) we show the scaling collapse (following Eq. 7 of main text) of the entanglement curves with $\tau_{c}=0.1$ and exponent $\nu=0.86$.
[…]
Results for $h=1$:
The numerical analysis of the survival probability and bipartite entanglement for $h=1$ is presented in panels (a) and (b) of Fig. S.6. In panel (a), we plot $\frac{dH}{d\tau}$ as a function of $\tau$ for system sizes $L=20,22,24,$ and $26$ at field strength $h=1$. For $L=26$, the transition point is found to be $\tau_{c}\approx 0.16$. Panel (b) displays a data collapse of the curves in the plot of $S(L/4,\tau)-S(L/4,\tau_{c})$ vs. $(\tau-\tau_{c})^{1/\nu}$ at the same field strength $h=1$. This collapse is obtained by considering system sizes $L=20,22,24,$ and $26$, yielding a critical exponent $\nu=1.331$ at $\tau_{c}=0.16$. These results indicate a clear transition from an area to a volume-law transition about $\tau_{c}\approx 0.16$.
Figure S.6: In panel (a), we have plotted $\frac{dH}{d\tau}$ vs. $\tau$ for $L=20,22,24$ and $26$ at the field strength $h=1$. At $h=1$, $\tau_{c}$ is found to be 0.16 approximately. In panel (b), we show the similar collapse of curves in the $S(L/4,\tau)-S(L/4,\tau_{c})$ vs. $(\tau-\tau_{c})^{1/\nu}$ plot for field strength $h=1$. This collapse is achieved by considering the system size of the spin chain $L=20,22,24$ and $26$. Here the value of the exponent is found to be $\nu=1.331$ at $\tau_{c}=0.16$.
@@
Figure S.7: In panel (a), we have plotted $\frac{dH}{d\tau}$ vs. $\tau$ for $L=100,200,300,500,1000$ at the field strength $h=\frac{3}{2}$. Even at $h=\frac{3}{2}$, $\tau_{c}$ also moves towards 0 as $L$ increases. In panel (b), we have plotted $\frac{dH}{d\sigma}$ vs. $\sigma$ at the same system sizes and field strength as FIG. 7(a), where $\sigma=\tau\sqrt{L}$. All the curves have a peak around $\sigma_{c}=1$.
Small size scaling of transition point:
We have already shown in the main manuscript that the transition point varies as $\tau_{c}=1/\sqrt{L}$. Here, we have examined how this transition behaves when considering only finite system sizes. In this regard, we have taken system sizes $L=16,18,20,22,24,26,$ and $28$. We have evaluated the scaling function for $h=\frac{1}{2}$ and $h=\frac{3}{2}$ by observing $|\tau_{c}-\tau_{c}^{\infty}|$ with system size $L$. Here, $\tau_{c}^{\infty}$ denotes the transition points at the thermodynamic limit, i.e., $L\rightarrow\infty$. The transition point at the thermodynamic limit for $h=\frac{1}{2}$ and $h=\frac{3}{2}$ is $\tau_{c}^{\infty}=0.107$ and, $\tau_{c}^{\infty}=0.103$ respectively. The scaling functions corresponding to $h=\frac{1}{2}$ and $h=\frac{3}{2}$ field strengths have been found to be $|\tau_{c}-\tau_{c}^{\infty}|=-0.109\ln{L}+0.416$ and $|\tau_{c}-\tau_{c}^{\infty}|=-0.081\ln{L}+0.317$, respectively. The least squared errors associated with the two fitting curves in Fig. 8(a) and Fig. 8(b) are given by $0.4405\%$ and $0.443\%$, respectively. Here it is to be noted that for any field strength the analytically achieved scaling function of the transition point, at large size limit, is proportional to $\frac{1}{\sqrt{L}}$, which does not match with the finite size scaling function for both the field strength.
Figure S.8: Here $\tau^{\infty}_{c}$ is the transition point at the thermodynamic limit, i.e., $L\rightarrow\infty$. In panel (a), we do finite size scaling of the transition point $\tau_{c}$ at field strength $h=\frac{1}{2}$. In panel (b), we do the finite size at $h=\frac{3}{2}$ scaling in the same way as in panel (a). In both panels, we use system size $L=16,18,20,22,24,26$, and $28$.
Figure S.9: Bipartite entanglement (of $L/4$: $3L/4$ of the system) just before (blue) and after (orange) the measurement as a function of time-step $n$ for $L=24$ and a) $\tau=0.06$ and b) $\tau=0.4$.
Figure S.10: (a) Entanglement (violet) and survival probability (orange) vs. $n\tau$ for $\tau=0.12$ and $L=24$. This is a typical trend observed for all $\tau$ values investigated. (b) Entanglement vs $n\tau$ is shown for several values of $\tau$. The first dip, which is actually the end point of the plateau, is marked by a black arrow.
Numerical Method of finding the exponent $\nu$: The detailed method of finding the exponent is $\nu$ in our work is as follows. Firstly, to find the data collapse, $n$ numbers of data files ($S$ vs $\tau$) corresponding to each system size $L$ are created. In our numerical analysis, we have created four data files, i.e $n=4$. For a given $\tau_{c}$ and $\nu$, we construct the data sets $S(\tau)-S(\tau_{c})$ vs $(\tau-\tau_{c})L^{1/\nu}$ for each L. If there is no data for $S$ for the given $\tau_{c}$, we use linear interpolation to find $S(\tau_{c})$. Now we select a window around $\tau_{c}$, say $\tau_{c}-\tau_{0}$ to $\tau_{c}+\tau_{0}$, and break it into a number of bins as per our raw data resolution. In each bin we have $S(\tau)-S(\tau_{c})$ values for different $L$. We calculate the mean of all the data in a particular bin and calculate the least squared error for each bin and calculate the total least squared error by summing over all the bins. Now we select the final $\tau_{c}$ and $\nu$ for which the total least squared error is minimum.
Difference in the role of measurement with the random circuit model:
As discussed in the main manuscript, in quantum circuits the transition originates from a competition between the measurement process, which suppresses entanglement, and the unitary evolution, which enhances it. In contrast, in our scenario the transition emerges even though the measurement can either increase or decrease the entanglement, depending on the specific case. This feature is depicted in Fig. S.10.
Dip of entanglement and the end point of the plateau: Here, we observe an interesting behavior of dip in entanglement in the dynamics and the endpoint of the plateau. In particular, the termination of the plateau region in the $R_{n}$ vs $n$ plot is accompanied by a sudden dip in entanglement dynamics. This simultaneous occurrence suggests that the existence of the plateau in the $R_{n}$ vs $n$ plot is closely linked to the behavior of the entanglement entropy. This feature is pictorially shown in Fig. S.10.